in a poll of 46 students 23 liked rap music 24 liked rock music,and 19 liked country music. of all the students 12 liked rap and country. 13 liked rap and rock, and 14 liked country and rock. of those students 9 liked all 3 types. how many students did not like any of these types.
Answer by stanbon(48558)

in a poll of 46 students 23 liked rap music 24 liked rock music,and 19 liked country music. of all the students 12 liked rap and country. 13 liked rap and rock, and 14 liked country and rock. of those students 9 liked all 3 types. how many students did not like any of these types.
-----------------
Draw three interlocking circles (Venn Diagram)
If you can fill in the various sections you will find the following number of students IN the circles:
# in Union = AUBUC-(A and B)-(A and C)-(B and C)+ (A and B and C)
= 23+24+19-12-13-14+9 = 36
The number not in the circles is 46-36= 10
================
Cheers,
Stan H.